# font=sf,footnotesize UN, Nantes]Nantes Université, CNRS, CEISAM UMR 6230, F-44000 Nantes, France

### v1

@@ 1. Introduction
On the theoretical side, several methods are available to get more accurate VTEs. 14 Firstly, one can further climb Jacob’s ladder of DFT, opting for double hybrids (DHs) rather than GHs or RSHs. 15, 16, 17, 18 Secondly, another approach involves employing the Bethe-Salpeter equation (BSE) formalism 19, 20 of many-body perturbation theory 21, 22 combined with the method. 23, 24, 25, 26 The so-called BSE/ approach advantageously washes out the XCF dependency, when a (partially) self-consistent scheme is used. 27, 28, 29 Thirdly, one can opt for correlated wave function approaches. Within this framework, the systematically improvable coupled-cluster (CC) route stands as a natural choice, offering a clear path for increases in accuracy with its CC2, 30, 31 CCSD, 32, 33, 34, 35, 36 CC3, 37, 38, 39 CCSDT, 40, 41, 42, 43, 44 CC4, 45, 46, 47, 48 CCSDTQ, 49, 50, 51, 45 and so forth declinations.
@@ 3.1 Theoretical Best Estimates
| Compound | State | TBE | Method | Compound | State | TBE | Method |
| Anthracene | () | 3.776 | Eq. 3 | (Ryd) | 3.169 | Eq. 4 | |
| () | 3.804 | Eq. 3 | (Ryd) | 3.693 | Eq. 4 | ||
| () | 5.022 | Eq. 3 | Heptazine | () | 2.717 | Eq. 2 | |
@@
To delve deeper into the accuracy of the TBEs, a valuable metric is , which defines the percentage of single-excitation character of the transition. Indeed, systematically increasing the degree of the CC excitation operator would not be the optimal strategy if doubly-excited states (characterized by low values of ) were involved. 158, 159 Among the 122 cases treated here, 62 show , indicating that they can be safely viewed as almost pure single excitations. Notably, this includes all 53 triplets, which comes as no surprise. 160, 52 Most of the remaining singlet states, i.e., 44, fall in the range, where CC3 (or CCSDT) generally performs extremely well. Indeed, we have recently demonstrated that corrections from quadruples (CC4 or CCSDTQ) are essentially negligible for such transitions, 48 although such observations can only be safely stated for smaller molecules. Next, 10 states are characterized by , for which both CC3 and CCSDT are likely no longer chemically accurate, but probably still superior to CASPT2 (or similar methods), based on our experience. 161, 162 Finally, Table 1 includes a few transitions with (6 ESs). In these cases, we expect the present TBEs to be too large, i.e., upper bounds of the exact VTEs, as corrections from quadruples typically downshift such states with a significant double excitation character. 48 For example, for the well-known dark ES of butadiene (), the CC3 and CCSDT errors, relative to the near-exact VTE, are +0.16 and +0.08 eV, respectively.52, 159
Another strategy to probe CC3’s accuracy is to investigate the differences between the CC3 and CCSDT values. Though this approach is not entirely foolproof, 48 it hints at the convergence of the CC family of methods. From the 26 singlet ESs for which we could compute this difference in the 6-31+G(d) basis set, it results a MAE of 0.034 eV which we regard as a positive indicator of our TBEs’ quality. Figure 2 shows the CC3-CCSDT VTE difference as a function of . As can be seen, there is no clean relationship between the two parameters, although the two largest deviations correspond to states with . Another interesting outcome, already noticed for substituted phenyls, 163 is that the CCSDT correction tends to be positive, meaning that the CC3 VTEs tend to be lower than their CCSDT counterparts for large molecules.
@@ 3.2 Wave Function Benchmarks
Having the above-described TBEs at hand, it is natural to assess the performances of lower-level wave function approaches. Indeed, the CC3/aug-cc-pVTZ VTEs listed in Table 1 are at the frontier of today’s technical feasibility. Thus, for practical applications, one would clearly wish to use methods that are computationally more affordable. The interested reader can find the full list of raw data in the SI for the 14 wave function methods tested here. In Table 2, we report the results of the statistical analysis: mean signed error (MSE), MAE, and standard deviation of the errors (SDE), together with the MAE obtained for various subgroups of ESs. In Figure 3, histograms representing the distribution of these errors are available. It should be recalled that: i) there are only 8 Rydberg transitions in our set; ii) CCSD(T)(a)⋆, CCSDR(3), and CCSDT-3 have been implemented for singlet ESs only; and iii) for some of the most expensive levels of theory [CCSDT-3 or ADC(3), for example], we were not able to obtain all possible VTEs despite our efforts.
| Full set | MAE of subsets | ||||||||
@@
For the present set of real-life dyes, CC2 emerges as particularly robust with a small MSE, an attractive SDE of 0.12 eV (indicating quite systematic errors) and MAE close to 0.10 eV for all ES families. Interestingly, based on QUEST, we found that the MAE of CC2 is 0.16 eV, 52 but this larger value was mostly driven by the small compounds. Indeed, the MAE is 0.11 eV for the 7–10 (non-H) atom molecules of QUEST, hinting that the convergence of the CC2 error with system size is likely reached. For comparisons, the MAE and SDE of CC2 were found to be 0.07 eV and 0.09 eV, respectively, in the study of Winter and coworkers, 77 who selected 0-0 experimental energies (for the lowest-lying ESs) as references. We believe that our statistical indicators are slightly higher due to the inclusion of more challenging (higher-energy) ESs in our set.
@@
As discussed in many works, 181, 77, 103, 78, 182, 179, 180, 52, 53, 163 ADC(2) does provide a rather similar accuracy as CC2. However, ADC(2) seems slightly less effective than its CC counterpart here, as evidenced by the higher values of the statistical indicators for most ES subgroups considered in Table 2, especially the transitions for which the advantage of CC2 over ADC(2) is clear. This finding differs from the one of Ref. 52 but parallels the outcomes of Winter and coworkers, 77 who reported a MAE of 0.08 eV and a SDE of 0.12 eV for ADC(2).
@@
Even though the overestimation tendency of CCSD is well-documented, 160, 183, 184, 185, 182, 186, 52, 53, 163 the MSE (0.18 eV) and MAE (0.20 eV) of Table 2 are surprisingly large, with an obvious unbalanced nature. For example, the CCSD error is larger for singlets than for triplets. It is also larger for than for transitions. What was even less expected is the large SDE of 0.16 eV, which clearly exceeds its CC2 counterpart of 0.12 eV. Indeed, the reverse SDE trend was noticed in QUEST, 52 CCSD delivering more systematic errors than CC2. In light of these outcomes, the interest in using CCSD rather than CC2 for applications on large compounds is definitely not justified. Adding perturbative corrections for the missing triple excitations helps, with SDE well below 0.10 eV for both CCSD(T)(a)⋆ and CCSDR(3). The MSEs of these two similar methods remain firmly positive, though. Unfortunately, this confirms 52, 163 that these two schemes become less effective when one considers larger molecules as they cannot downshift the CCSD VTEs enough. To solve this issue, one can turn to iterative triples with CCSDT-3, but this expensive method shares with CC3 the unpleasant scaling with system size. Note that CCSDT-3 still overestimates our TBEs, albeit by a small acceptable margin.
@@
As originally reported by us, 104 ADC(3) does not improve upon ADC(2), while ADC(2.5), where one simply average the ADC(2) and ADC(3) VTEs, 52, 105, 163 appears again to be a relatively affordable approach providing accurate results. All statistical errors of ADC(2.5) are close to 0.06 eV, this method impressively outperforming both CCSD(T)(a)⋆ and CCSDR(3) on the present set, with an absolute accuracy comparable to the one of CCSDT-3, but with an underestimation trend.
@@
Briefly, the most effective spin-scaled approach evaluated here is likely SOS-ADC(2) with the QC parameters, 101 which delivers an accuracy comparable to the one of CC2, for the same formal scaling but with a smaller prefactor. SCS-CC2 (with the TM parameters) yields a very small SDE but at the cost of a clear overshooting trend, a conclusion paralleling previous studies. 75, 78, 187, 188, 52, 53, 163 Both the SOS-ADC(2) and SOS-CC2 with TM parameters appear as less reliable options for the current set.
@@ 3.3.1 Gauge Invariance Correction
As mentioned in Section 2, one can include gauge invariance corrections in TD-DFT calculations involving (hybrid) meta-GGA XCFs. The impact of these corrections was carefully studied and quantified in previous works. 136, 137, 138 As enforcing the TDA breaks gauge invariance, it has been argued that applying the gauge corrections at the TDA-DFT level is not necessarily well-grounded. 137 Nevertheless, we have found that the magnitude of the corrections obtained with TDA-DFT and “full” TD-DFT for a given transition are extremely similar (see Figures S1 and S2 in the SI).
@@ 3.3.3 Effect of the TDA
As expected, 196, 151, 86, 197, 198, 195, 199, 137 applying the TDA induces an almost systematic increase in the transition energies, a statement holding true for both TD-DFT and BSE/. The few exceptions are related to states characterized by a significant orbital mixing with an admixture modified by the application of the TDA. B97X-2 seems particularly prone to deliver negative TDA corrections. We present histograms of the TDA corrections for B3LYP, SOGGA-11X, M06-SX, CAM-B3LYP, B97, cLH20t, as well as two BSE/ schemes in Figure 6. The graphs for the other XCFs as well as Tables of statistical data can be found in Sections S4.2 and S5.2 of the SI.
@@ 3.3.4 TD-DFT vs TBE
We now turn towards the assessment of all XCFs against the TBEs. Full data can be found in Table S24-S25 (for TD-DFT) and S28 (for BSE) in the SI, together with histograms of the errors, whereas a graphical representation of the MSEs, MAEs, and SDEs can be found in Figure 8. We start our analysis below by concentrating on the TD-DFT data before turning our attention to the BSE/ results.
@@
As expected, for the GHs, the evolution of the TD-DFT MSEs (roughly) follows the percentage of EXX: 13 it is negative for XCFs with low EXX percentage and becomes close to zero from 40% of EXX. The impact of applying the TDA and thus upshifting all values is also clear. For example, it decreases the MSE of PBE0 from to eV while increasing its M06-2X counterpart from +0.03 to +0.12 eV. Within the TDA, all RSHs have positive MSEs, except for tCAM-B3LYP and mCAM-B3LYP, whereas in TD-DFT, B97X-D shows a tiny average deviation of eV. Two of the tested LHs, namely, cLH12ct-SsirPW92 and cLH20t, also show small negative (positive) MSEs with TD-DFT (TDA-DFT). Amongst the tested DHs, three of them (PBE0-DH, SOS-B88PP86, and SOS-PBEPP86) show absolute MSEs smaller than 0.10 eV, irrespective of the application of the TDA.
@@
Let us now turn to the MAEs. The MAEs for each ES subsets can be found in Tables S24 and S25 of the SI. For the most accurate XCFs, considering only transitions with a dominant single-excitation character () slightly decreases the MAEs, which is the expected trend since adiabatic TD-DFT is not equipped to describe properly ESs having a partial double-excitation character. In contrast, the relative deviations obtained for , , and Rydberg ESs is strongly dependent on the XCF, with a magnified XCF-sensitivity for the latter type of states. 137 Amongst the GHs, the smallest MAEs are obtained for TD-cM06-2X and TD-M06-2X, the former functional delivering an error as small as 0.14 eV for the transitions showing , a rather remarkable result. It is noteworthy that, with these two XCFs, the final error is driven by the singlet (and not triplet) ESs. Comparisons for smaller molecules using Thiel’s set have indeed highlithed TD-M06-2X as a valuable method for singlet-triplet transitions. 201 When applying the TDA, the smallest GH MAE is obtained with BMK: 0.19 eV or 0.16 eV for the transitions. Amongst the RSHs, the smallest MAEs are 0.21 eV for TD-B97M-V and 0.19 eV with TDA-(c)M06-SX and TDA-CAM-B3LYP. Globally, the long-range corrected XCFs with large attenuation parameters (B97 and LC-PBE) are significantly less satisfying, but the present set does not contain clear CT ESs that might require such XCFs. 72 Amongst the three tested LHs, cLH20t appears to be the most accurate for the present dataset with MAEs of 0.19 eV (TD-DFT) and 0.16 eV (TDA-DFT). The accuracy of TD-DFT/TDA-DFT can be further improved by using the more computationally demanding DHs with MAEs of approximately 0.15 eV for PBE0-DH and PBE-QIDH. Even smaller values are obtained with XCFs designed to reproduce transition energies, namely SOS-B88PP86 and SOS-PBEPP86, the later providing a quite astonishing MAE of 0.11 eV when combined with the TDA. However, relatively poor estimates are obtained for the few Rydberg transitions considered here.
@@
An important feature is consistency and this is measured by the SDE which assesses the systematic character of the errors. Gratifyingly, in both TD-DFT and TDA-DFT, the smallest SDEs are obtained with cM06-2X in the GH series and cLH20t in the LH series: the MAE and SDE performances parallel each other for these two groups of XCFs. This is not the case for the RSHs at the TD-DFT level, since tCAM-B3LYP, which was tuned to reproduce transition energies in diarylethene photochromes, delivers a small SDE of 0.22 eV, despite a rather large MAE of 0.35 eV. Within the TDA, the smallest (RSH) SDEs are reached with mCAM-B3LYP, CAM-B3LYP, B97X-D, and B97M-V. In the DH series, PBE-QIDH provides a SDE of 0.12 eV at the TD-DFT level (a set that does not contain triplets, however). When enforcing the TDA, a SDE of 0.15 eV can be obtained with the same functional, as well as with the already mentioned SOS-B88PP86 and SOS-PBEPP86 XCFs.

### v2

@@ 1. Introduction
On the theoretical side, several methods are available to get more accurate VTEs. 14 Firstly, one can further climb Jacob’s ladder of DFT, opting for double hybrids (DHs) rather than GHs or RSHs. 15, 16, 17, 18 Secondly, another approach involves employing the Bethe-Salpeter equation (BSE) formalism 19, 20 of many-body perturbation theory 21, 22 combined with the method. 23, 24, 25, 26 The so-called BSE/ approach advantageously washes out the XCF dependency, when a (partially) self-consistent scheme is used. 27, 28, 29 The diagonalization of the electron-hole BSE Hamiltonian in the (de)excitation space offers the same computational scaling as Casida’s formulation of TD-DFT, 3 namely to get iteratively the lowest eigenvalues. Similarly, the preceding calculations, required to get the quasiparticle energies and screened Coulomb potential that enter the BSE Hamiltonian, can be performed with scaling using resolution-of-the-identity (RI) techniques. Thirdly, one can opt for correlated wave function approaches. Within this framework, the systematically improvable coupled-cluster (CC) route stands as a natural choice, offering a clear path for increases in accuracy with its CC2, 30, 31 CCSD, 32, 33, 34, 35, 36 CC3, 37, 38, 39 CCSDT, 40, 41, 42, 43, 44 CC4, 45, 46, 47, 48 CCSDTQ, 49, 50, 51, 45 and so forth declinations.
@@ 3.1 Theoretical Best Estimates
ine Compound | State | TBE | Method | Compound | State | TBE | Method |
ine Anthracene | () | 3.757 | Eq. 1 | (Ryd) | 3.169 | Eq. 4 | |
() | 3.782 | Eq. 1 | (Ryd) | 3.693 | Eq. 4 | ||
() | 5.012 | Eq. 1 | Heptazine | () | 2.717 | Eq. 2 | |
@@ 3.1.1 On the accuracy of the TBEs
To delve deeper into the accuracy of the TBEs, a valuable metric is , which defines the percentage of single-excitation character of the transition. Indeed, systematically increasing the degree of the CC excitation operator would not be the optimal strategy if doubly-excited states (characterized by low values of ) were involved. 158, 159 Among the 122 cases treated here, 62 show , indicating that they can be safely viewed as almost pure single excitations. Notably, this includes all 53 triplets, which comes as no surprise. 160, 52 Most of the remaining singlet states, i.e., 44, fall in the range. In other words, 106-out-of-122 TBEs are associated with . For these transitions, both CC3 and CCSDT generally performs extremely well. Indeed, for such transitions we demonstrated, on the one hand, that corrections from quadruples (CC4 or CCSDTQ) are essentially negligible, 48 whereas, on the other hand, second-order multireference methods, e.g. CASPT2 and NEVPT2, are significantly less accurate than CC3 with errors of approximatively 0.10-0.13 and 0.02 eV, respectively.161 Next, 10 states are characterized by , for which both CC3 and CCSDT are likely no longer chemically accurate, but probably still superior to CASPT2 (or similar methods), based on our experience. 161, 162 Finally, Table 1 includes a few transitions with (6 ESs). In these cases, we expect the present TBEs to be too large, i.e., upper bounds of the exact VTEs, as corrections from quadruples typically downshift such states with a significant double excitation character. 48 For such transitions in rather large compounds, it is difficult to know if the CC3/CCSDT or CASPT2/NEVPT2 would be the most accurate. We can only refer to smaller structures for which indisputable TBEs are at hand. 52, 159 One can obtain hints from smaller compounds. For the well-known dark ES of butadiene (), the CC3 and CCSDT errors, relative to the near-exact VTE, are +0.15 and +0.08 eV, respectively, whereas the CASPT2 and NEVPT2 deviations are +0.22 and +0.18 eV, respectively. 52, 159 For the lowest ES of benzene (), the CC3, CASPT2 and NEVPT2 errors are respectively +0.18, +0.12, and +0.32 eV.159 Consequently, there is no clear indication that multireference values would be more accurate than the present CC-based TBEs for our set of data.
Another strategy to probe CC3’s accuracy is to investigate the differences between the CC3 and CCSDT values. Though this approach is not entirely foolproof in assessing the impact of further (quadruple) CC corrections, 48 it hints at the convergence of the CC family of methods. From the 29 singlet ESs for which we could compute this difference in the 6-31+G(d) basis set, it results a MAE of 0.034 eV which we regard as a positive indicator of our TBEs’ quality. We note that the use of this small basis set should not be an issue. Indeed, we have shown that Eq. 2 provides excellent estimates of the actual CCSDT/AVTZ results, with a MAE of 0.005 eV, i.e., the CC3-CCSDT differences seems reasonably insensitive to the basis set size.48 Figure 2 shows these differences as a function of . As can be seen, there is no clean relationship between the two parameters, although the two largest deviations correspond to states with . Another interesting outcome, already noticed for substituted phenyls, 163 is that the CCSDT correction tends to be positive, meaning that the CC3 VTEs tend to be lower than their CCSDT counterparts for large molecules.
3.1.2 Comparisons with literature data
@@ 3.2 Wave Function Benchmarks
Having the above-described TBEs at hand, it is natural to assess the performances of lower-level wave function approaches. Indeed, the CC3/aug-cc-pVTZ VTEs listed in Table 1 are at the frontier of today’s technical feasibility. Thus, for practical applications, one would clearly wish to use methods that are computationally more affordable. The interested reader can find the full list of raw data in the SI for the 14 wave function methods tested here. In Table 2, we report the results of the statistical analysis: mean signed error (MSE), MAE, and standard deviation of the errors (SDE), together with the MAE obtained for various subgroups of ESs. In Figures 3, S1, and S2, histograms representing the distribution of these errors are available. It should be recalled that: i) there are only 8 Rydberg transitions in our set; ii) CCSD(T)(a)⋆, CCSDR(3), and CCSDT-3 have been implemented for singlet ESs only; and iii) for some of the most expensive levels of theory [CCSDT-3 for instance], we were not able to obtain all possible VTEs despite our efforts.
ine | Full set | MAE of subsets | |||||||
@@
For the present set of real-life dyes, CC2 emerges as particularly robust with a small MSE, an attractive SDE of 0.12 eV (indicating quite systematic errors) and MAE close to 0.10 eV for all ES families. Interestingly, based on QUEST, we found that the MAE of CC2 is 0.16 eV, 52 but this larger value was mostly driven by the small compounds. Indeed, the MAE is 0.11 eV for the 7–10 (non-H) atom molecules of QUEST, hinting that the convergence of the CC2 error with system size is likely reached. Nevertheless, one notes a bimodal distribution of the CC2 errors in Figure 3. As can be seen in Figure S1, this originates from the tendency of CC2 to underestimate (overestimate) the VTE of both Rydberg and () ESs, which might cause issues if a balanced description of all state types is required. For comparisons, the MAE and SDE of CC2 were found to be 0.07 eV and 0.09 eV, respectively, in the study of Winter and coworkers, 77 who selected 0-0 experimental energies (for the lowest-lying ESs) as references. We believe that our statistical indicators are slightly higher due to the inclusion of more challenging (higher-energy) ESs in our set.
@@
As discussed in many works, 181, 77, 103, 78, 182, 179, 180, 52, 53, 163 ADC(2) does provide a rather similar accuracy as CC2. However, ADC(2) seems slightly less effective than its CC counterpart here, as evidenced by the higher values of the statistical indicators for most ES subgroups considered in Table 2, especially the transitions for which the advantage of CC2 over ADC(2) is clear, since ADC(2) more significantly underestimates the VTEs of those transitions than CC2 (See Figures S1 and S2). This finding differs from the one of Ref. 52 but parallels the outcomes of Winter and coworkers, 77 who reported a MAE of 0.08 eV and a SDE of 0.12 eV for ADC(2).
@@
Even though the overestimation tendency of CCSD is well-documented, 160, 183, 184, 185, 182, 186, 52, 53, 163 the MSE (0.18 eV) and MAE (0.20 eV) of Table 2 are surprisingly large, with an obvious unbalanced nature. In particular, the CCSD positive biais is markedly larger for singlets than for triplets (Figure S1). It is also larger for than for transitions, though in that case the error pattern seem better than their CC2 counterparts. What was even less expected is the large SDE of 0.16 eV, which clearly exceeds its CC2 counterpart of 0.12 eV. Indeed, the reverse SDE trend was noticed in QUEST, 52 CCSD delivering more systematic errors than CC2. In light of these outcomes, the interest in using CCSD rather than CC2 for applications on large compounds is difficult to justify in general. Adding perturbative corrections for the missing triple excitations helps, with SDE well below 0.10 eV for both CCSD(T)(a)⋆ and CCSDR(3). The MSEs of these two similar methods remain firmly positive, though. Unfortunately, this confirms 52, 163 that these two schemes become less effective when one considers larger molecules as they cannot downshift the CCSD VTEs enough. To solve this issue, one can turn to iterative triples with CCSDT-3, but this expensive method shares with CC3 the unpleasant scaling with system size. Note that CCSDT-3 still overestimates our TBEs, albeit by a small acceptable margin.
@@
As originally reported by us, 104 ADC(3) does not improve upon ADC(2), while ADC(2.5), where one simply average the ADC(2) and ADC(3) VTEs, 52, 105, 163 appears again to be a relatively affordable approach providing accurate results. Figure S2 clearly shows that the error patterns of ADC(2) and ADC(3) are mirrored for the different types of ESs, explaning the success of ADC(2.5). Indeed, all statistical errors of ADC(2.5- are close to 0.06 eV and it impressively outperforms both CCSD(T)(a)⋆ and CCSDR(3) on the present set, with an absolute accuracy comparable to the one of CCSDT-3, but with an underestimation trend.
@@
Briefly, the most effective spin-scaled approach evaluated here is likely SOS-ADC(2) with the QC parameters, 101 which delivers an accuracy comparable to the one of CC2, for the same formal scaling but with a smaller prefactor. SCS-CC2 (with the TM parameters) yields a very small SDE but at the cost of a clear overshooting trend, a conclusion paralleling previous studies. 75, 78, 187, 188, 52, 53, 163 Both the SOS-ADC(2) and SOS-CC2 with TM parameters appear as less reliable options for the current set. All tested spin-scaled methods also provide reasonably balanced error patterns for singlets and triplets (Figure S2), which is likely advantageous when estimating singlet-triplet gaps.
@@ 3.3.1 Gauge Invariance Correction
As mentioned in Section 2, one can include gauge invariance corrections in TD-DFT calculations involving (hybrid) meta-GGA XCFs. The impact of these corrections was carefully studied and quantified in previous works. 136, 137, 138 As enforcing the TDA breaks gauge invariance, it has been argued that applying the gauge corrections at the TDA-DFT level is not necessarily well-grounded. 137 Nevertheless, we have found that the magnitude of the corrections obtained with TDA-DFT and “full” TD-DFT for a given transition are extremely similar (see Figures S3 and S4 in the SI).
@@ 3.3.3 Effect of the TDA
As expected, 196, 151, 86, 197, 198, 195, 199, 137 applying the TDA induces an almost systematic increase in the transition energies, a statement holding true for both TD-DFT and BSE/. The few exceptions are related to states characterized by a significant orbital mixing with an admixture modified by the application of the TDA. B97X-2 seems particularly prone to deliver negative TDA corrections. We present histograms of the TDA corrections for B3LYP, SOGGA-11X, M06-SX, CAM-B3LYP, B97, cLH20t, as well as two BSE/ schemes in Figure 6. The graphs for the other XCFs as well as Tables of statistical data can be found in Sections S5.2 and S6.2 of the SI.
@@ 3.3.4 TD-DFT vs TBE
We now turn towards the assessment of all XCFs against the TBEs. Full data can be found in Table S24-S25 (for TD-DFT) and S29 (for BSE) in the SI, together with histograms of the errors, whereas a graphical representation of the MSEs, MAEs, and SDEs can be found in Figure 8. We start our analysis below by concentrating on the TD-DFT data before turning our attention to the BSE/ results.
@@
As expected, for the GHs, the evolution of the TD-DFT MSEs (roughly) follows the percentage of EXX: 13 it is negative for XCFs with low EXX percentage and becomes close to zero from 40% of EXX. The impact of applying the TDA and thus upshifting all values is also clear. For example, it decreases the MSE of PBE0 from to eV while increasing its M06-2X counterpart from to eV. Within the TDA, all RSHs have positive MSEs, except for tCAM-B3LYP and mCAM-B3LYP, whereas in TD-DFT, B97X-D shows a tiny average deviation of eV. Two of the tested LHs, namely, cLH12ct-SsirPW92 and cLH20t, also show small negative (positive) MSEs with TD-DFT (TDA-DFT). Amongst the tested DHs, three of them (PBE0-DH, SOS-B88PP86, and SOS-PBEPP86) show absolute MSEs smaller than 0.10 eV, irrespective of the application of the TDA.
@@
Let us now turn to the MAEs. The MAEs for each ES subsets can be found in Tables S24 and S25 of the SI. For the most accurate XCFs, considering only transitions with a dominant single-excitation character () slightly decreases the MAEs, which is the expected trend since adiabatic TD-DFT is not equipped to describe properly ESs having a partial double-excitation character. In contrast, the relative deviations obtained for , , and Rydberg ESs is strongly dependent on the XCF, with a magnified XCF-sensitivity for the latter type of states. 137 Amongst the GHs, the smallest MAEs are obtained for TD-cM06-2X and TD-M06-2X, the former functional delivering an error as small as 0.14 eV for the transitions showing , a rather remarkable result. It is noteworthy that, with these two XCFs, the final error is driven by the singlet (and not triplet) ESs. Comparisons for smaller molecules using Thiel’s set have indeed highlithed TD-M06-2X as a valuable method for singlet-triplet transitions. 201 When applying the TDA, the smallest GH MAE is obtained with BMK: 0.19 eV or 0.16 eV for the transitions. Amongst the RSHs, the smallest MAEs are 0.21 eV for TD-B97M-V and 0.19 eV with TDA-(c)M06-SX and TDA-CAM-B3LYP. Globally, the long-range corrected XCFs with large attenuation parameters (B97 and LC-PBE) are significantly less satisfying, but the present set does not contain clear CT ESs that might require such XCFs. 72 Amongst the three tested LHs, cLH20t appears to be the most accurate for the present dataset with MAEs of 0.19 eV (TD-DFT) and 0.15 eV (TDA-DFT). The accuracy of TD-DFT/TDA-DFT can be also improved by using the more computationally demanding DHs with, e.g., MAEs of approximately 0.15 eV for PBE0-DH and PBE-QIDH. Even smaller values are obtained with XCFs designed to reproduce transition energies, namely SOS-B88PP86 and SOS-PBEPP86, the later providing a quite astonishing MAE of 0.10 eV when combined with the TDA. However, relatively poor estimates are obtained for the few Rydberg transitions considered here.
@@
An important feature is consistency and this is measured by the SDE which assesses the systematic character of the errors. Gratifyingly, in both TD-DFT and TDA-DFT, the smallest SDEs are obtained with cM06-2X in the GH series and cLH20t in the LH series: the MAE and SDE performances parallel each other for these two groups of XCFs. This is not the case for the RSHs at the TD-DFT level, since tCAM-B3LYP, which was tuned to reproduce transition energies in diarylethene photochromes, delivers a small SDE of 0.22 eV, despite a rather large MAE of 0.35 eV. Within the TDA, the smallest (RSH) SDEs are reached with mCAM-B3LYP, CAM-B3LYP and B97X-D. In the DH series, PBE-QIDH provides a SDE of 0.12 eV at the TD-DFT level (a set that does not contain triplets, however). When enforcing the TDA, a SDE of 0.15 eV can be obtained with the same functional, as well as with the already mentioned SOS-B88PP86 and SOS-PBEPP86 XCFs.